OpenStudy (anonymous):

find the inverse of the matrix if it exists p=[4,4 4,5]

OpenStudy (across):

\[\begin{vmatrix} 4 & 4\\ 4 & 5 \end{vmatrix}=20-16=4\neq0.\]It has an inverse.

OpenStudy (mathmate):

The inverse of a 2x2 matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] is \[\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] divided by the determinant of the matrix (ac-bd).

OpenStudy (mathmate):

Sorry, the determinant should read (ad-bc).

OpenStudy (across):

Given \(A\) is a \(2\times2\) matrix, it's inverse can be computed by: \[A^{-1}=\frac{1}{ad-cb}\begin{bmatrix} d & -b\\ -c & a \end{bmatrix}.\]Do you know how to do that, nickyd63?

OpenStudy (anonymous):

no that's what i need help with i just do not understand it

OpenStudy (across):

All you have to do is plug the matrix values into the above expression, like this:\[A^{-1}=\frac{1}{4\cdot5-4\cdot4}\begin{bmatrix} 5 & -4\\ -4 & 4 \end{bmatrix}.\]Can you simplify that?

OpenStudy (anonymous):

No, im still trying to figure out how you did that

OpenStudy (mathmate):

From Across's matrix, you'll see that the top left element a has exchanged places with element d in the inverse, and the elements b and c have changed sign. That's how she got the inverse. \[(1/4)\left[\begin{matrix}5 & -4 \\ -4 & 4\end{matrix}\right]\] However, you still have to divide each element by the determinant, calculated using ad-bc, or in the given case, (4)(5)-(4)(4)=4. The final result becomes: \[\left[\begin{matrix}5/4 & -1 \\ -1 & 1\end{matrix}\right]\]